# Coherent polarization self-rotation

### v1

@@ V Discussion
CPSR introduces a robust and versatile framework for quantum optics and spectroscopy in highly broadened, optically thick spin ensembles. It operates effectively in the spin-exchange relaxation-free regime, under high buffer gas pressure and large optical depth, and exhibits narrow-band, near-unity coherent absorption or strong gain, even under conditions that typically hinder other coherent schemes. The performance of CPSR is governed by the spin-light coupling rate $\Omega$ [as given by Eq. (5) for weak probe absorption]. Maximal signal absorption at $\omega\approx\omega_{\mathrm{a}}$, corresponding to unity contrast $\mathcal{C}=1$, occurs when $\Omega=2\gamma$, where $\gamma$ is the spin decoherence rate. For larger values of $\Omega$, the spin-light coupling strength and signal amplification at $\omega\approx-\omega_{\mathrm{a}}$ increase, with contrast reaching $|\mathcal{C}|=\Omega/(2\gamma)$.
@@
The coupling can be enhanced by increasing the control power $P_{\mathrm{c}}$ up to some saturation point, since both $\Omega\propto P_{\mathrm{c}}$ and $\gamma\propto P_{\mathrm{c}}$ at high powers. In our system, the contrast $|\mathcal{C}|$ would saturate at approximately 40 for $P_{\mathrm{c}}>50$ mW with $30$ GHz pump detuning. Further enhancement may be achieved by increasing the optical pumping rate and raising the cell temperature, provided that the pumping power is sufficient to avoid depletion of the pumping beam. If depletion becomes significant, spatial effects, such as inhomogeneous light shifts, must also be considered. Miniaturized vapor cells (e.g., 10 mm3) may offer a promising route toward compact, high-performance CPSR applications. In rubidium, under feasible conditions (e.g., 10 mW laser power, 250 C cell temperature, and 1000 Torr buffer gas), CPSR could achieve $|\mathcal{C}|>7$ with a linewidth on the order of $500$ Hz.
Furthermore, CPSR can serve as a platform for quantum information applications, where non-coherent schemes such as PSR are inapplicable. As shown in Appendix B.4, Eqs. (1) and (2) remain valid when quantum operators replace the corresponding classical variables. Although the basic CPSR mechanism couples only a single optical quadrature to one spin quadrature, it can be extended to a full (dual-quadrature) quantum interface by implementing a double-pass geometry [32], with the present work indicating it can work with buffer gas and in the SERF regime. Such an interface could enable a variety of quantum operations, including efficient storage and retrieval of nonclassical states of light in optically thick alkali ensembles.
A key motivation for developing the CPSR scheme is to enable quantum optical coupling to long-lived nuclear spins of noble gases, mediated by dense alkali vapors. Previous demonstrations in this domain have typically relied on weak, adiabatic excitation of the alkali spins to resolve the millihertz-scale spin resonances of the noble-gas nuclei [22]. In contrast, CPSR permits substantial excitation of the alkali spins, opening the door to the strong-coupling regime of alkali-noble-gas interactions [21]. This regime could facilitate storage with bandwidths that are orders of magnitude larger than those achievable in the adiabatic-coupling limit [20, 33].
@@ A.2 Effect of optical pumping
Figure 6: Signal amplification, given by $S_{2}^{\text{out}}e^{d(\Delta)}/S_{2}^{\text{in}}-1$ at the anti-resonance of CPSR, as a function of the optical pumping rate $R_{\mathrm{pump}}/q$. The value of $R_{\mathrm{pump}}/q$ is independently extracted from time-domain measurements of the spin decoherence rate. As optical pumping increases both spin polarization and decoherence, the amplification exhibits a maximum, observed when $R_{\mathrm{pump}}/q=27.5$ Hz is comparable to the intrinsic spin decoherence rate $\gamma(R_{\mathrm{pump}}=0)=21.5$ Hz. The solid line is a result of the calibrated numerical model.
@@ B.1 Detailed model
The detailed model is based on Bloch equations, incorporating the relevant physical processes illustrated in Fig. 1c [14, 35]. The excited-state population is adiabatically eliminated, assuming a predominant ground-state population. Frequent spin-exchange collisions among alkali atoms drive this population towards a spin-temperature distribution, characterized by the spin polarization $p_{\mathrm{a}}$ and the slowing-down factor $q$ [29, 14, 28]. These same collisions are also responsible for the emergence of spin-exchange relaxation-free (SERF) behavior.
@@
$+\frac{2}{q}\frac{\Delta}{\Gamma}\sigma(\Delta)s_{3}\hat{\mathbf{x}}\times\mathbf{f}+n_{\mathrm{a}}(\frac{R_{\mathrm{pump}}}{2}\hat{\mathbf{z}}+\sigma(\Delta)s_{3}\hat{\mathbf{x}}),$
@@
For naturally abundant rubidium, we treat the mean-field spin as an effective composite spin comprising 72.2% ${}^{85}\mathrm{Rb}$ and 27.8% ${}^{87}\mathrm{Rb}$. This treatment is justified by the rapid spin-exchange collisions between all rubidium atoms in the regime $R_{\mathrm{se}}\gg\omega_{\mathrm{a}},\omega$, which equilibrate the spin temperature (and thus the spin polarization $p_{\mathrm{a}}$) across both species. As a result, the effective slowing-down factor is given by $q(p_{\mathrm{a}})=0.722q_{I=5/2}(p_{\mathrm{a}})+0.278q_{I=3/2}(p_{\mathrm{a}})$, ranging between $10.85.44=q(1)<q(p_{\mathrm{a}})<q(0)=$ 10.8, and the system can be described by a single coherent spin field $\mathbf{f}$. For potassium, $q=q_{I=3/2}$ since the nuclear spin of the naturally abundant composition is almost entirely $I=3/2$.
@@
$A$ Beam cross-sectional area, equal to the cell cross-section. The cell’s volume is $A\ell$. $\mathrm{cm^{2}}$
@@
$p_{\mathrm{a}}$ Alkali-spin polarization $-1\leq p_{\mathrm{a}}\leq 1$. unitless
@@
$q$ Slowing-down factor for the alkali spins (assuming a spin-temperature distribution), $q=0.278q_{I=3/2}+0.722q_{I=5/2}$ with $q_{I=3/2}=\frac{6+2p_{\mathrm{a}}^{2}}{1+p_{\mathrm{a}}^{2}}$ and $q_{I=5/2}=\frac{38+52p_{\mathrm{a}}^{2}+6p_{\mathrm{a}}^{4}}{3+10p_{\mathrm{a}}^{2}+3p_{\mathrm{a}}^{4}}$ for natural isotopic abundant rubidium, and $q=q_{I=3/2}$ for potassium [37, 38]. unitless
@@ B.2 Simplified model and analytical solution
[…]
@@
$+n_{\mathrm{a}}(\frac{R_{\mathrm{pump}}}{2}\hat{\mathbf{z}}+\frac{1-e^{-d(\Delta)}}{d(\Delta)}\sigma(\Delta)s_{3}^{\text{in}}\hat{\mathbf{x}}),$ (7a)
@@ B.3 Numerical solution
The theoretical curves shown in Figs. 2, 3, 4, and 6 are obtained by numerically solving the coupled equations of the detailed model, Eqs. (7). The simulations assume a spatially uniform optical pumping rate $R_{\mathrm{pump}}$ and Larmor frequency $\omega_{\mathrm{a}}$, as well as a probe beam that uniformly illuminates the entire cross-sectional area of the cell. For rubidium, natural isotopic abundance is used. The parameters used in these calculations are listed in Table 2.
For the calculations of the rubidium experiments in Figs. 2, 3, and 6, the model parameters yield a bare spin decoherence rate of $\gamma(R_{\mathrm{pump}}=\omega_{\mathrm{a}}=0)=13.2$ Hz. The discrepancy with the experimentally measured value (21.5 Hz) may arise from spatial inhomogeneities in the pumping rate and lightshifts, caused by partial absorption of the pumping beam.
Parameter Calculations for Figs. 2, 3 Calculation for Fig. 6 Calculation for Fig. 4
$\ell$ $1\,\mathrm{cm}$ $1\,\mathrm{cm}$ (effective for a 14 mm diameter sphere)
$A$ $1\,\mathrm{cm^{2}}$ $1.5\,\mathrm{cm^{2}}$
$\Gamma$ $2.6\,\mathrm{GHz}$ $12\,\mathrm{GHz}$
$\Delta_{\text{pump}}$ $45\,\mathrm{GHz}$ $100\,\mathrm{GHz}$
$P_{\mathrm{c}}$ $15.2\,\mathrm{mW}$ $25\,\mathrm{mW}$
$P_{\mathrm{s}}$ $20\,\mathrm{\mu W}$ $35\,\mathrm{\mu W}$
$P_{\text{pump}}$ $60\,\mathrm{mW}$ $0-100\,\mathrm{mW}$ $1.6\,\mathrm{mW}$ (effective over the cell)
$\omega_{\mathrm{a}}$ $268\,\mathrm{Hz}$ (Fig. 2) $290-400\,\mathrm{Hz}$ $29\,\mathrm{Hz}$
$T$ $154\,\mathrm{Celsius}$ $185\,\mathrm{Celsius}$
$\Delta$ $89\,\mathrm{GHz}$ $130\,\mathrm{GHz}$ $215\,\mathrm{GHz}$
$R_{\mathrm{sd}}$ $98\,\mathrm{Hz}$ $94\,\mathrm{Hz}$ $27\,\mathrm{Hz}$
@@ B.4 Simplified Heisenberg-Langevin model for quantum operators
where $\hat{a}_{z}(x)$ and $\hat{a}_{y}(x)$ are the annihilation operators for the $z$-polarized control and $y$-polarized signal modes, respectively, at spatial position $x$. These operators satisfy the bosonic commutation relations $[\hat{a}_{\chi}(x),\hat{a}^{\dagger}_{\chi^{\prime}}(x^{\prime})]=c\delta_{\chi,\chi^{\prime}}\delta(x-x^{\prime})$, and $\langle\hat{a}^{\dagger}_{\chi}(x)\hat{a}_{\chi}(x)\rangle$ is the local photon flux of the mode $\chi\in\{y,z\}$. The classical Stokes parameters correspond to the expectation values of the quantum Stokes operators.
Under the Holstein-Primakoff approximation for the optical field [31], the control mode is treated as a strong, classical field, while the signal mode remains a weak quantum field. This allows the quantum operators associated with the control to be replaced by classical variables. In particular, the operator $\hat{s}_{1}$ is replaced by its expectation value $\hat{s}_{1}\mapsto\langle\hat{s}_{1}\rangle=s_{1}$. Similarly, we apply the Holstein-Primakoff approximation to the spin density operators $\hat{f}_{x}$, $\hat{f}_{y}$, and $\hat{f}_{z}$. Due to the optical pumping, $\hat{f}_{z}$ can be treated as a macroscopic, classical number $\hat{f}_{z}\mapsto\langle\hat{f}_{z}\rangle=f_{z}$. Under these approximation, the commutation relations for the light and spin operators become $[\hat{s}_{2}(x),\hat{s}_{3}(x^{\prime})]=ics_{1}(x)\delta(x-x^{\prime})$ and $[\hat{f}_{x}(\mathbf{r}),\hat{f}_{y}(\mathbf{r}^{\prime})]=if_{z}(\mathbf{r})\delta(\mathbf{r}-\mathbf{r}^{\prime})$.

### v2

@@ V Discussion
CPSR introduces a robust, versatile framework for quantum optics and spectroscopy in highly broadened, optically thick spin ensembles. It operates effectively in the spin-exchange relaxation-free regime under high buffer-gas pressure and large optical depth, exhibiting narrow-band, near-unity coherent absorption or strong gain even under conditions that typically hinder other coherent schemes. Performance is governed by the spin-light coupling rate $\Omega$ [given by Eq. (5) for weak probe absorption]. Maximal absorption at $\omega\approx\omega_{\mathrm{a}}$, corresponding to unity contrast $\mathcal{C}=1$, occurs for $\Omega=2\gamma$, where $\gamma$ is the spin decoherence rate. For larger $\Omega$, the coupling strength and amplification at $\omega\approx-\omega_{\mathrm{a}}$ increase, with contrast reaching $|\mathcal{C}|=\Omega/(2\gamma)$.
@@
The coupling can be increased by raising the control power $P_{\mathrm{c}}$ up to saturation, occurring since at high power both $\Omega\propto P_{\mathrm{c}}$ and $\gamma\propto P_{\mathrm{c}}$. In our system, $|\mathcal{C}|$ would saturate near $\sim 40$ for $P_{\mathrm{c}}>50$ mW at $30$ GHz pump detuning. Further enhancement is possible by increasing the optical pumping rate and the cell temperature, provided that the pump power is sufficient to avoid depletion. If depletion becomes significant, spatial effects such as inhomogeneous light shifts must also be accounted for. Miniaturized vapor cells (e.g., 10 mm3) offer a route to compact, high-performance CPSR applications. In rubidium, under feasible conditions (e.g., 10 mW laser power, 250 C cell temperature, and 1000 Torr buffer gas pressure), CPSR could reach $|\mathcal{C}|>7$ with a linewidth on the order of $500$ Hz.
While CPSR benefits from the long coherence times and high optical depth of the SERF regime and exhibits magnetically sensitive spectral features, it differs conceptually from conventional SERF magnetometers. In SERF magnetometry, spin orientation is commonly excited magnetically and read out optically via Faraday rotation. In CPSR, both excitation and readout are optical, driven by vector lightshifts, forming a two-photon interface that generates the observed optical spectra. Consequently, SERF magnetometers typically operate in the weakly coupled probe regime, where the spin coherence is minimally perturbed by the probe, whereas CPSR relies on strong, bidirectional spin-light coupling. Optical driving also enables spatial resolution set by the beam size rather than global magnetic excitation, albeit with reduced sensitivity at small spatial scales due to atomic thermal motion.
CPSR relies on the dispersive light-matter interaction, which preserves coherence between the optical field and the collective spin. This differs from PSR, in which spin orientation is generated by dissipative optical pumping via the circular component of the elliptically polarized beam, and the resulting polarization rotation arises from dispersive Faraday rotation. The dissipative optical absorption in PSR limits the realization of a fully coherent light-matter interface. Accordingly, CPSR requires an auxiliary pump to establish spin orientation. Its coupling strength increases with the on-resonance optical depth $d(0)$ $\Omega\propto d(0)$ in Eq. (5)], whereas PSR exhibits an optimal optical depth near $d(0)\approx 1$, with performance degrading as $d(0)$ increases.
Notably, since CPSR is both coherent and optically driven, it realizes the Faraday Interaction (Quantum-Non-Demolition, QND) Hamiltonian [30]. Optical fields can carry nonclassical states, thereby enabling a quantum interface with the collective spin and opening avenues for quantum information applications. In Appendix B.4, we show that the input-output relations extend to quantum operators [Eq. (B.4)]. CPSR enables nonclassical protocols previously demonstrated in spin-preserving-coated cells [31] to be implemented in buffer-gas SERF cells. The key distinction is the presence of rapid spin-exchange interactions and strong optical broadening. In earlier demonstrations, atomic interactions were considered a source of decoherence and mostly neglected, focusing on the collective coupling to the optical field [32]. In contrast, SERF enables long coherence times and high optical depth without specialized coatings.
We emphasize that CPSR is not intended to outperform state-of-the-art SERF magnetometers in magnetic-field sensitivity. These have demonstrated sensitivity at the spin projection noise limit and QND readout [33, 34], and their exceptional sensitivity has been achieved through careful optimization of experimental parameters, such as increasing optical depth [2]. CPSR, on the other hand, enables the use of collective spin orientation as a quantum resource for quantum information and precision measurement applications.
As near-term milestones, CPSR in the nonclassical regime enables (i) generation of nonlocal entanglement between two spatially separated ensembles via QND measurement of a joint spin quadrature [35], and (ii) a double-pass geometry that implements a spin-light beam-splitter Hamiltonian for efficient storage and retrieval of weak optical fields [36], with extensions toward on-demand single-photon generation [37].
These milestones open a path to quantum-enhanced metrology in the acoustic band. LIGO is a particularly significant test case: it is one of the few platforms where quantum enhancement has already delivered system-level sensitivity gains. It operates in the acoustic band, where the SERF regime combines high optical depth with suppression of spin-exchange-induced decoherence, and back-action-evasive readout via QND (negative mass reference frame) has been proposed [38]. Recent advances, including fast, sub-cycle squeezing [39], and wavelength transduction of squeezing [40], outline a viable path to acoustic band implementations [41], for which CPSR in the SERF regime is well suited.
A practical consideration for such protocols is spin polarization. Moderate nonclassical performance typically requires $\gtrsim 70\%$ polarization [19], whereas $95\%$ is readily achievable in alkali vapors [42]. Pump-induced relaxation can be mitigated by shutting the pump off during the quantum sequence.
Finally, CPSR provides fast, on-resonance quantum-optical coupling to long-lived nuclear spins of noble gases, via a dense alkali vapor. Whereas prior work addressed noble-gas nuclei via weak, adiabatic alkali excitation to resolve mHz resonances [22], CPSR supports substantial coherent alkali excitation, unlocking the strong-coupling regime [21]. Strong coupling could facilitate orders-of-magnitude improvement of light storage bandwidth [20, 43].
@@ A.2 Effect of optical pumping
Figure 6: Signal amplification, given by $S_{2}^{\text{out}}e^{d(\Delta)}/S_{2}^{\text{in}}-1$ at the anti-resonance of CPSR, as a function of the optical pumping rate $R_{\mathrm{pump}}/q$. The value of $R_{\mathrm{pump}}/q$ is independently extracted from time-domain measurements of the spin decoherence rate. As optical pumping increases both spin polarization and decoherence, the amplification exhibits a maximum, observed when $R_{\mathrm{pump}}/q=27.5$ Hz is comparable to the intrinsic spin decoherence rate $\gamma(R_{\mathrm{pump}}=0)=21.5$ Hz. The solid purple line is the prediction of the simplified model, and the dashed black line is that of the calibrated numerical model.
@@ B.1 Detailed model
The detailed model is based on Bloch equations, incorporating the relevant physical processes illustrated in Fig. 1[14, 45]. The excited-state population is adiabatically eliminated, assuming a predominant ground-state population. Frequent spin-exchange collisions among alkali atoms drive this population towards a spin-temperature distribution, characterized by the spin polarization $p_{\mathrm{a}}$ and the slowing-down factor $q$ [29, 14, 28]. These same collisions are also responsible for the emergence of spin-exchange relaxation-free (SERF) behavior.
@@
[…]
@@
For naturally abundant rubidium, we treat the mean-field spin as an effective composite spin comprising 72.2% $^{85}\mathrm{Rb}$ and 27.8% $^{87}\mathrm{Rb}$. This treatment is justified by the rapid spin-exchange collisions between all rubidium atoms in the regime $R_{\mathrm{se}}\gg\omega_{\mathrm{a}},\omega$, which equilibrate the spin temperature (and thus the spin polarization $p_{\mathrm{a}}$) across both species. As a result, the effective slowing-down factor is given by $q(p_{\mathrm{a}})=0.722q_{I=5/2}(p_{\mathrm{a}})+0.278q_{I=3/2}(p_{\mathrm{a}})$, ranging between $5.44=q(1)<q(p_{\mathrm{a}})<q(0)=10.8$, and the system can be described by a single coherent spin field $\mathbf{f}$. For potassium, $q=q_{I=3/2}$ since the nuclear spin of the naturally abundant composition is almost entirely $I=3/2$.
@@
AA Beam cross-sectional area, equal to the cell cross-section. The cell’s volume is $A\ell$. $\mathrm{cm^{2}}$
@@
$p_{\mathrm{a}}$ Alkali-spin polarization $1\leq p_{\mathrm{a}}\leq 1$. unitless
@@
$q$ Slowing-down factor for the alkali spins (assuming a spin-temperature distribution), $q=0.278q_{I=3/2}+0.722q_{I=5/2}$ with $q_{I=3/2}=\frac{6+2p_{\mathrm{a}}^{2}}{1+p_{\mathrm{a}}^{2}}$ and $q_{I=5/2}=\frac{38+52p_{\mathrm{a}}^{2}+6p_{\mathrm{a}}^{4}}{3+10p_{\mathrm{a}}^{2}+3p_{\mathrm{a}}^{4}}$ for naturally abundant rubidium, and $q=q_{I=3/2}$ for potassium [47, 48]. unitless
@@ B.2 Simplified model and analytical solution
The essence of CPSR is captured by a simplified model, derived by assuming uniformity of the pumping rate, Larmor frequency, and probe beam intensity over the cell volume, as well as negligible optical pumping by the probe, as detailed below. First, we assume that both the optical pumping rate $R_{\mathrm{pump}}$ and the Larmor frequency $\omega_{\mathrm{a}}$ are spatially homogeneous across the vapor cell. Second, we assume that the probe beam uniformly covers the cell’s cross-section and that its attenuation along the propagation axis is weak, i.e. the off-resonant optical depth satisfies $d(\Delta)=\ell n_{\mathrm{a}}\sigma(\Delta)\ll 1$. Under these conditions, we directly solve Eq. (6e) for $s_{3}$, obtaining $s_{3}(x,t)=e^{-d(\Delta)x/\ell}s_{3}^{\text{in}}(t)$, compute its spatial average $\frac{1}{\ell}\int_{0}^{\ell}s_{3}(x,t)dx=\frac{1-e^{-d(\Delta)}}{d(\Delta)}s_{3}^{\text{in}}$, and substitute this into Eq. (6a). As a result, Eqs. (6a) reduce to purely temporal equations, and the spin field $\mathbf{f}$ becomes spatially uniform. Solving the resulting system for $s_{1,2}$, we obtain the following dynamical input-output relations:
@@
[…]
@@ B.3 Numerical solution
The theoretical curves shown in Figs. 2, 3, 4, and 6. For the simplified model [Eq. (4)], we derive $\Omega,\omega_{\mathrm{a}},\gamma$ from the model’s parameters. For the numerical model, solutions are obtained by numerically solving Eqs. (7). The simulations assume a spatially uniform optical pumping rate $R_{\mathrm{pump}}$ and Larmor frequency $\omega_{\mathrm{a}}$, as well as a probe beam that uniformly illuminates the entire cross-sectional area of the cell. For rubidium, natural isotopic abundance is used. The parameters used in these calculations are listed in Table 2.
For the calculations of the rubidium experiments in Figs. 2, 3, and 6, the model parameters yield a bare spin decoherence rate of $\gamma(R_{\mathrm{pump}}=\omega_{\mathrm{a}}=0)=13.2$ Hz. The discrepancy with the experimentally measured value (21.5 Hz) may arise from spatial inhomogeneities in the pumping rate and lightshifts, caused by partial depletion of the pumping beam.
Parameter Calculations for Calculation for Calculation for
Figs. 2, 3 Fig. 6 Fig. 4
$\ell$ 1 cm 1 cm (effective for a 14 mm diameter sphere)
AA 1 cm2 1.5 cm2
$\Gamma$ 2.6 GHz 12 GHz
$\Delta_{\text{pump}}$ 45 GHz 100 GHz
$P_{\mathrm{c}}$ 15.2 mW 25 mW
[…]
$P_{\text{pump}}$ 60 mW 0–100 mW 1.6 mW (effective over the cell)
$\omega_{\mathrm{a}}$ 268 Hz (Fig. 2) 290–400 Hz 29 Hz
$T$ 154 C 185 C
$\Delta$ 89 GHz 116 GHz 215 GHz
$R_{\mathrm{sd}}$ 98 Hz 94 Hz 27 Hz
@@ B.4 Simplified Heisenberg-Langevin model for quantum operators
where $\hat{a}_{z}(x)$ and $\hat{a}_{y}(x)$ are the annihilation operators for the $z$-polarized control and $y$-polarized signal modes, respectively, at spatial position $x$. These operators satisfy the bosonic commutation relations $\hat{a}_{\chi}(x),\hat{a}^{\dagger}_{\chi^{\prime}}(x^{\prime})]=c\delta_{\chi,\chi^{\prime}}\delta(x-x^{\prime})$, and $\langle\hat{a}^{\dagger}_{\chi}(x)\hat{a}_{\chi}(x)\rangle$ is the local photon flux of the mode $\chi\in\{y,z\}$. The classical Stokes parameters correspond to the expectation values of the quantum Stokes operators.
Under the Holstein-Primakoff approximation for the optical field [31], the control mode is treated as a strong, classical field, while the signal mode remains a weak quantum field. This allows the quantum operators associated with the control to be replaced by classical variables. In particular, the operator $\hat{s}_{1}$ is replaced by its expectation value $\hat{s}_{1}\mapsto\langle\hat{s}_{1}\rangle=s_{1}$. Similarly, we apply the Holstein-Primakoff approximation to the spin density operators $\hat{f}_{x}$, $\hat{f}_{y}$, and $\hat{f}_{z}$. Due to the optical pumping, $\hat{f}_{z}$ can be treated as a macroscopic, classical number $\hat{f}_{z}\mapsto\langle\hat{f}_{z}\rangle=f_{z}$. Under these approximation, the commutation relations for the light and spin operators become $\hat{s}_{2}(x),\hat{s}_{3}(x^{\prime})]=ics_{1}(x)\delta(x-x^{\prime})$ and $\hat{f}_{x}(\mathbf{r}),\hat{f}_{y}(\mathbf{r}^{\prime})]=if_{z}(\mathbf{r})\delta(\mathbf{r}-\mathbf{r}^{\prime})$.